OpenStudy (anonymous):

An original data set has the values 11, 8, 7, 12, and 7. What would need to be added to each value to change the mean to ? possible answers: 3 6 7 8

OpenStudy (anonymous):

Please someone help I need to get this done or I don'g get to watch Doctor Who for a week

OpenStudy (anonymous):

@MDoodler Can you help?

OpenStudy (anonymous):

Oh you are changing the mean to 15 Dieu, I'm so stupid

OpenStudy (anonymous):

I just need to know what the answer is

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

You just have a question mark at the end. What are we changing the means to?

OpenStudy (anonymous):

The mean right now is \[ \frac{7+7+8+11+12}{5} = 35.4\]

OpenStudy (anonymous):

If you add 3 to each value \[ \frac{10+10+11+14+15}{5} = 48 \]